A system at equilibrium is represented by the following chemical equation: H₂(g) + I₂(g) <--> 2HI(g) + 9.40 kJ At equilibrium, this system has a [H₂] = [I₂] = 0.666 M and [HI] = 1.234 M. Calculate the value of K for this system.
1
Expert's answer
2020-06-10T07:41:06-0400

K = [HI]2/([H2]*[I2]) = 1.2342/(0.666*0.666) = 3.43